# On the Convergence of Federated Learning Algorithms without Data Similarity

### v1

@@ V-C Discussions
From Figures 1 and 2, under the same fixed step size, FedAvg outperforms FedProx in full-precision and error-feedback updates regarding solution accuracy and convergence speed, especially when the data similarity is high. For the IID case, at $K=100$, FedAvg attains $80\%$ test accuracy while FedProx achieves $65\%$ test accuracy. This occurs since, given the small learning rate $0.1$, the regularization term of the proximal update in FedProx dominates the whole objective function; thus, the next local iterate $x_{i}^{k}$ is almost the same as the current global iterate $x^{k}$. Also, error-feedback algorithms perform worse than their full-precision counterparts, particularly when the data similarity is low. For the Non-IID1 case, at $K=200$, error-feedback algorithms obtain $40\%$ test accuracy while full-precision algorithms reach $70\%$ test accuracy. This happens because the top-$k$ sparsifier in error-feedback algorithms, unlike full-precision algorithms, introduces biased information.
Turning to Figures 345 and 6, similar to the results with fixed step sizes, FedAvg tends to outperform FedProx; error-feedback algorithms are inclined to have a worse performance than their full-precision counterparts when the step sizes are diminishing and step-decay.
@@ VI Conclusions
We have presented a unified analysis framework for federated algorithms on non-convex problems without data similarity assumptions. This framework requires the worst-case convergence bound in (1) and applies convergence theorems that incorporate fixed, diminishing, and step-decay step size schedules. We illustrate how to apply this framework to establish strong convergence results for full-precision and error-feedback federated algorithms with step sizes independent of data similarity parameters under standard conditions on objective functions. Finally, numerical experiments on training the CNN models on the MNIST and FashionMNIST datasets demonstrate the performance of these federated algorithms under various data similarity conditions.
@@ Appendix F Proof of Proposition 1
where $\beta^{k}={\alpha^{k}}/{2}-L(\alpha^{k})^{2}$, $A_{1}^{k}=\mathbf{E}\left\|\nabla f(x^{k})-\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\nabla f_{i}(x_{i}^{k,t})\right\|^{2}$, and also $A_{2}^{k}=\mathbf{E}\left\|\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}B_{i}^{k,t}\right\|^{2}$. If $\alpha^{k}\leq 1/(\sqrt{6}L)$, then $\alpha^{k}\leq 1/(2L)$. Hence, $\beta^{k}\geq 0$ and

### v2

@@ V-C Discussions
V-C1 (Error-feedback) FedAvg vs. (Error-feedback) FedProx Algorithm
Figures  4 and  5 show that with the same fixed step size, FedAvg surpasses FedProx in both full-precision and error-feedback updates in terms of solution accuracy and convergence speed, particularly when data similarity is high. For instance, in the IID case at $K=100$, FedAvg achieves an 80% test accuracy, whereas FedProx only reaches 65%. This disparity arises because, with a small learning rate of 0.1, the proximal update’s regularization term in FedProx becomes dominant, causing the next local iterate $x_{i}^{k}$ to be nearly identical to the current global iterate $x^{k}$.
V-C2 Error-Feedback vs. Full-Precision Federated Learning Algorithms
We also observed that error-feedback algorithms generally underperform compared to their full-precision counterparts, especially when data similarity is low. For example, in the Non-IID1 scenario at $K=200$, error-feedback algorithms achieve a 40% test accuracy, whereas full-precision algorithms attain 70%. This performance gap is due to the top-$k$ sparsifier in error-feedback algorithms introducing biased information, unlike in full-precision algorithms.
V-C3 Effect of Different Step Size Regimes
Figures  6,  7,  8, and  9 demonstrate consistent trends with the fixed step size results. Similar to those results, FedAvg tends to outperform FedProx, and error-feedback algorithms generally exhibit poorer performance than their full-precision counterparts under diminishing and step-decay step size regimes. Furthermore, our theoretical findings are validated, showing that these different algorithms can converge without requiring step sizes to be coupled to data similarity.
@@ VI Conclusions
In this work, we have introduced a unified analysis framework for federated algorithms on non-convex problems without relying on data similarity assumptions. This framework employs the worst-case convergence bound in the general non-negative system  (1) and utilizes convergence theorems that incorporate fixed, diminishing, and step-decay step size schedules. We demonstrated how to apply this framework to achieve strong convergence results for both full-precision and error-feedback federated algorithms. This includes FedAvg, FedProx, error-feedback FedAvg, and error-feedback FedProx, all with step sizes that are independent of data similarity parameters under standard conditions on objective functions. Finally, we substantiated our theoretical findings with numerical experiments, training CNN models on the MNIST and FashionMNIST datasets. These experiments showcase the performance of these federated algorithms under various data similarity conditions.
@@ Appendix F Proof of Proposition 1
where $\beta^{k}={\alpha^{k}}/{2}-L(\alpha^{k})^{2}$, $A_{1}^{k}=\mathbf{E}\left\|\nabla f(x^{k})-\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\nabla f_{i}(x_{i}^{k,t})\right\|^{2}$, and also $A_{2}^{k}=\mathbf{E}\left\|\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}B_{i}^{k,t}\right\|^{2}$.
If $\alpha^{k}\leq 1/(\sqrt{6}L)$, then $\alpha^{k}\leq 1/(2L)$. Hence, $\beta^{k}\geq 0$ and